OpenStudy (anonymous):

What is the pH of a solution with a hydrogen-ion concentration of 4.2×10^-10 M

OpenStudy (nikato):

ph=-log[H+] Where [H+] is your concentration of H+

OpenStudy (anonymous):

Can you walk me through an example?

OpenStudy (nikato):

Do you have an example you want to work on?

OpenStudy (anonymous):

Not an particular one. Any is fine.

OpenStudy (nikato):

Well. If you have some problems on pH, we can work on it together. Or else I'll have to make up one right now and the numbers might not be too pretty, but will use same process

OpenStudy (anonymous):

Yeah I'm looking in my book I'll post one in a few.

OpenStudy (nikato):

Ok

OpenStudy (anonymous):

Here's one What is the pH of a solution with a hydrogen-ion concentration of \[4.2\times10^{-10}M\]

OpenStudy (anonymous):

So I know my known is \[[H^{+}]=4.2\times 10^{-10}\] and our equation is \[pH=-\log[H^{+}]\] And my unkown is pH=?

OpenStudy (anonymous):

So what do I do next?

OpenStudy (nikato):

Substitution Plug in 4.2x10-10 in [H+] And use calculator to solve

OpenStudy (nikato):

If you did it correct, you should get around 9.38

OpenStudy (anonymous):

So \[pH=-\log[4.2\times 10^{-10}]\] \[\log[H^{+}]=\log[4.2 \times 10^{-10}]\] \[=-9.37675\] \[pH=-(-9.38)\] \[pH=9.38\] And it's basic

OpenStudy (nikato):

100% correct!